OpenStudy (happykiddo):

Given: x – (5 – 3x + 2) = 15 + 3x Prove: x = 22 What are the missing justification steps, in order? Commutative Property of Addition, Combine Like Terms, Distributive Property Associative Property of Addition, Distributive Property, Multiplication Property of Equality Associative Property of Addition, Commutative Property of Addition, Distributive Property Associative Property of Addition, Distributive Property, Addition Property of Equality

OpenStudy (happykiddo):

oh darn the chart didnt come out right just a sec

OpenStudy (anonymous):

x – (5 – 3x + 2) = 15 + 3x x -5 + 3x - 2 = 15 + 3x 4x - 7 = 15 + 3x 4x - 3x = 15 + 7 x = 22

OpenStudy (happykiddo):

the ones with the question marks under are the ones that have the unknown property and this would be were the answers come in

OpenStudy (anonymous):

I prove it right?

OpenStudy (happykiddo):

yes

OpenStudy (happykiddo):

Mathematical Statement Justification x – (5 – 3x + 2) = 15 + 3x Given x – (5 + 2– 3x) = 15 + 3x ????? x – (7– 3x) = 15 + 3x ???? x – 7 + 3x = 15 + 3x ???? x + 3x – 7 = 15 + 3x Commutative Property of Addition 4x – 7 = 15 + 3x Combine Like Terms x – 7 = 15 Subtraction Property of Equality x = 22 Addition Property of Equality

OpenStudy (happykiddo):

use these instead above

OpenStudy (anonymous):

I really don't understand it

OpenStudy (happykiddo):

oh well its fine i just wanted to check my answer

OpenStudy (happykiddo):

i think its the first one

OpenStudy (anonymous):

you just want to ckeck? I thought it's a multiple choices

OpenStudy (happykiddo):

it is the options were above